OpenStudy (anonymous):

First I'd like to say I'm not a kid trying to get you to solve my homework. I've been out of scholl now for 20 years and its been a long time since I've done this stuff so be nice please. Can I solve a 2 variable equation without having a set of equations. The actual problem is 120=20x + 4y?

OpenStudy (dumbcow):

yes and no .... yes, the equation can be solved for a specific variable "ex: solve for y" means get "y" by itself on one side of equation no, you can't solve a single equation for a solution (a specific point) , because technically every point on the line would be a solution

OpenStudy (pratyush5):

i think you can. yu just need to subtitue the value of y in terms of x

OpenStudy (pratyush5):

substitute*

OpenStudy (anonymous):

If you have two variables you need two equations to solve .. The same goes for a 3 variable equation: you need 3 equations.

OpenStudy (pratyush5):

i am in doubt

OpenStudy (anonymous):

this works for anything . If you have two missing varialbes in a single equation like surface area , you need someway to relate one of those variables with what you're given

OpenStudy (chihiroasleaf):

yes you can.., you just need to find the value of x and y that satisfies the equation.. since you just have one equation, then you'll have lots of solution for the equation...,

OpenStudy (anonymous):

Well here's the thing I already know the answer x=5.77 and y=1.15 but how do you solve for x and y?

OpenStudy (anonymous):

The original equation again was 120=20x+4y and as I said x=5.77 and y=1.15

OpenStudy (dumbcow):

in that case, you need a 2nd equation

OpenStudy (chihiroasleaf):

x = 5.77 and y = 1.15 is one of the solutions if you have only one equation.., how do you get these number? :)

OpenStudy (anonymous):

Thats why I'm asking dont know?

OpenStudy (anonymous):

I only found answer because I kept plugging in numbers til I found it. I had a diagram that gave this info. There is a torque of 120 inlbs on a 12 in base plate. I'm trying to solve for tension per bolt on the base there are 4 bolts and the bolts are 2 inches and 10 inches from the pivot. So to translate this into and equation you have 120= (x*10"*2bolts)+(y*2"*2bolts) there was a diagram that showed the 5.77 and 1.15 as the tensions per bolt and the explanation said a simple calculation can be done to find these answers?

OpenStudy (whpalmer4):

Are you sure there aren't any other relationships between those numbers? Where are you reading this?

OpenStudy (whpalmer4):

It is worth noting that the the ratio between those two numbers is very close to 5, which is also the ratio between the bolt distances from the pivot.

OpenStudy (whpalmer4):

It's just the equation for a lever arm in action. As Archimedes said, give me a lever long enough and a fulcrum on which to place it and I'll move the Earth. So you had two equations to solve: \[120=20x + 4y\]\[x = (10/2)y = 5y\]As you have one already in terms of the other, I'd do substitution:\[120 = 20(5y) + 4y = 104y \rightarrow y = 120/104 = 15/13 \approx 1.15\]\[x =5(15/13) = 75/13 \approx 5.77\]

OpenStudy (anonymous):

whpalmer.. sorry i logged off a bit ago. how did you arrive with the second equation? The x=(10/2)y=5y don't understand? and Thank you very much ahead of time

OpenStudy (anonymous):

Also Im new to this site if i knew how I'd give you a medal

OpenStudy (whpalmer4):

The lever formula is just the force = the lever arm * weight. Think of a see-saw. If the pivot (fulcrum) is at r = 0, you've got one force pushing at r = 2" and another at r = 10". The force due to the one at 10" is (10/2) times the one at 2" because the lever arm is longer. Same idea as putting a piece of pipe over a wrench handle when you've got a bolt that won't move.

OpenStudy (anonymous):

Thanks you are very smart and it's greatly appreciated. I never thought of thinking in those terms. You definitely got some good answers.